# Some new cases of Zilber–Pink in $Y(1)^{3}$

### v1

@@ 2.6. G-functions associated with an abelian scheme
Consider the situation in Section 2.4 and let $\pi\colon\mathcal{A}\to C$ be an abelian scheme of relative dimension $g$, also defined over $K$. We write $H^{1}_{DR}(\mathcal{A}/C)$ for the relative algebraic de Rham cohomology, equipped with its Gauss–Manin connection
@@
By [Del70, Prop. 6.14], the $\mathcal{O}_{C}$-sheaf $H^{1}_{DR}(\mathcal{A}/C)$ is locally-free. Suppose that it is free. Equivalently, we can choose a basis $\{\omega_{1},\ldots,\omega_{2g}\}$ of global sections. Now let $v\in\Sigma_{K}$ and let $\Delta_{v}$ be an open subspace of $C^{v\textrm{-}\mathrm{an}}$ containing $s^{v\textrm{-}\mathrm{an}}_{0}$ such that $H^{1}_{DR}(\mathcal{A}/C)^{v\textrm{-}\mathrm{an}}(\Delta_{v})$ possesses a basis of horizontal sections $\{\gamma_{v,1},\ldots,\gamma_{v,2g}\}$. (For $v\in\Sigma_{K,\infty}$, the existence of $\Delta_{v}$ is classical—any simply connected open subset would suffice. For $v\in\Sigma_{K,f}$, the existence of $\Delta_{v}$ is guaranteed by [Ked22, Prop. 9.3.3].) Let
@@
Lemma 8.
Define $Y_{v}:=\Omega_{v}\cdot\Omega_{v}(s^{v\textrm{-}\mathrm{an}}_{0})^{-1}$ and denote its entries by $Y_{v,mn}$ (that is, $1\leq m,n\leq 2g$). Let $F_{v,mn}$ denote the Taylor series of $Y_{v,mn}$ with respect to $x$. Then the $F_{v,mn}$ are G-functions, belonging to $K[[X]]$, independent of $v\in\Sigma_{K}$.
@@
It is straightforward to verify that the matrix $F_{v}:=(F_{v,mn})_{mn}$ satisfies a linear differential equation $\frac{d}{dx}(-)=A(-)$ induced from the action of $\nabla(\frac{d}{dx})$ on $H^{1}_{DR}(\mathcal{A}/C)$, where $A\in\mathrm{M}_{2g}(K((X)))$. Starting from the fact that $F_{v}(0)=I_{2g}$ has entries in $K$, it follows that $F_{v,nm}\in K[[X]]$ for all $m,n\in\{1,\ldots,2g\}$. By [And89, p3, Thm. B], therefore, $F_{v,nm}$ is a G-function for all $m,n\in\{1,\ldots,2g\}$. Furthermore, $F_{v}$ is the unique solution of $\frac{d}{dx}(-)=A(-)$ in $\mathrm{M}_{2g}(K((X)))$ satisfying $F_{v}(0)=I_{2g}$, so it is independent of $v$. ∎
By virtue of Lemma 8, we denote $F_{v,mn}$ by $F_{mn}$ and we refer to the $F_{mn}$ for $1\leq m,n\leq 2g$ as the G-functions associated with $\mathcal{A}\to C$, $x$, and $\{\omega_{1},\ldots,\omega_{2g}\}$.
@@ 2.7. Relations
Let $F_{1},\dotsc,F_{n}\in\overline{\mathbb{Q}}[[X]]$. We say that a homogeneous polynomial $\tilde{Q}\in\overline{\mathbb{Q}}[X][X_{1},\dotsc,X_{n}]$ is a functional relation between $F_{1},\dotsc,F_{n}$ if
@@
Let $\xi\in\overline{\mathbb{Q}}$ and let $Q\in\overline{\mathbb{Q}}[X_{1},\dotsc,X_{n}]$ be a homogeneous polynomial. Let $K$ be a number field which contains $\xi$ and all of the coefficients of $F_{1},\dotsc,F_{n}$ and $Q$. We use the following definitions.
@@
    $Q$ is a trivial relation between $F_{1},\dotsc,F_{n}$ at $\xi$ if it is the specialisation at $X=\xi$ of a functional relation $\tilde{Q}\in\overline{\mathbb{Q}}[X][X_{1},\dotsc,X_{n}]$ between $F_{1},\dotsc,F_{n}$, where $\tilde{Q}$ is homogeneous of the same degree as $Q$.
@@ 2.8. Simple neighbourhoods
Return to the situation of Section 2.4 and let $\mathfrak{C}$ be a regular integral flat $\mathcal{O}_{K}$-scheme with an isomorphism $\mathfrak{C}_{K}\to C$. Let $v\in\Sigma_{K,f}$ and let $\mathfrak{C}_{v{\dagger}}$ denote the formal completion of $\mathfrak{C}$ along the fibre $\mathfrak{C}_{k_{v}}$. Let $\mathfrak{C}_{v{\dagger}}^{\mathrm{rig}}$ denote the associated rigid analytic space (see [Ber96, 0.2.6]). Then there is a canonical open immersion $\mathfrak{C}_{v{\dagger}}^{\mathrm{rig}}\to C^{v\textrm{-}\mathrm{an}}$ [Ber96, Prop. 0.3.5]. We say that an open subspace $U$ of $C^{v\textrm{-}\mathrm{an}}$ is simple if $U\subset\mathfrak{C}_{v{\dagger}}^{\mathrm{rig}}$ and the reduction map $\operatorname{red}_{v}:\mathfrak{C}_{v{\dagger}}^{\mathrm{rig}}\to\mathfrak{C}_{k_{v}}$ is constant on $U$.
The following is standard.
Lemma 9.
Let $U$ be a simple open subspace of $C^{v\textrm{-}\mathrm{an}}$ and let $t\in\mathfrak{C}(k_{v})$ denote the (constant) value of $\operatorname{red}_{v}$ on $U$. Then $\operatorname{red}_{v}^{-1}(t)$ is equal to the completion $\mathfrak{C}_{v,\mathrm{for}{\dagger}}$ of $\mathfrak{C}$ along $t$. If $t$ is a smooth point of $\mathfrak{C}_{k_{v}}$, then $\operatorname{red}_{v}^{-1}(t)$ is isomorphic to the rigid open unit disc.
@@ 3. Relations for elliptic schemes
In this section, we construct $v$-adic relations between the G-functions associated with elliptic schemes having isogenous fibres over a fixed base point. In Section 3.1, we consider a pair of elliptic schemes. In Section 3.3, we introduce a third elliptic scheme, in order to obtain a single relation for all places of good ordinary reduction of the fixed fibres.
@@ 3.1. Isogenous pairs
Let $C^{\prime}$ be a smooth projective geometrically irreducible algebraic curve, defined over a number field $K$, and let $\mathfrak{C}^{\prime}$ be a regular $\mathcal{O}_{K}$-model of $C^{\prime}$ (see [Liu06, Definition 10.1.1] for the definition; such a model exists by [Liu06, Cor. 8.3.45]).
@@
[…]
For $i\in\{1,2\}$, let $\mathfrak{G}^{\prime}_{i}\to\mathfrak{C}^{\prime}$ be a semiabelian scheme such that for some non-empty Zariski open subset $C\subset C^{\prime}$ the base-change $\mathcal{E}_{i}:=\mathfrak{G}^{\prime}_{i}\times_{\mathfrak{C}^{\prime}}C$ is an elliptic scheme over $C$. For $i\in\{1,2\}$, let $\mathfrak{C}\subset\mathfrak{C}^{\prime}$ denote a non-empty Zariski open subscheme such that $\mathfrak{G}_{i}:=\mathfrak{G}^{\prime}_{i}\times_{\mathfrak{C}^{\prime}}\mathfrak{C}$ is an elliptic scheme over $\mathfrak{C}$ (such a subscheme exists because the toric rank is upper semi-continuous [Lan13, Lemma 3.3.1.4]). Suppose that $\mathfrak{C}_{K}=C$.
Let $s_{0}\in C(K)$ and suppose that there exists a $K$-isogeny $f_{0}:\mathcal{E}_{1,s_{0}}\to\mathcal{E}_{2,s_{0}}$ of degree $N_{0}:=\deg(f_{0})$. We will say that $v\in\Sigma_{K,f}$ is a place of good (resp. bad) reduction if one (equivalently, all) of the $\mathcal{E}_{i,s_{0}}$ have good (resp. bad) reduction at $v$.
Choose bases $\{\omega_{i},\eta_{i}\}$ for $H^{1}_{DR}(\mathcal{E}_{i}/C)$ such that
(1) $\omega_{1,s_{0}}=f^{*}_{0}\omega_{2,s_{0}}\text{ and }\eta_{1,s_{0}}=f^{*}_{0}\eta_{2,s_{0}},$
where
@@
denotes the pullback induced by $f_{0}$. Let $x\in K(C)$ be a local parameter at $s_{0}$ and let $F_{imn}\in K[[X]]$ denote the G-functions associated with $\mathcal{E}_{i}\to C$, $x$, and $\{\omega_{i},\eta_{i}\}$. Let
$\mathcal{G}:=\{F_{imn}:1\leq i,m,n\leq 2\},$
and, for each $v\in\Sigma_{K}$, define
$R_{v}:=\min\{1,R(F^{v\textrm{-}\mathrm{an}}):F\in\mathcal{G}\}.$
For each $v\in\Sigma_{K,f}$, let $r_{v}$ and $U_{v}$ denote, respectively, the positive real number and the open subspace of $C^{v\textrm{-}\mathrm{an}}$ afforded to us by Proposition 20. For each $v\in\Sigma_{K,\infty}$, let $U_{v}$ denote an open subspace of $C^{v\textrm{-}\mathrm{an}}$ such that $x^{v\textrm{-}\mathrm{an}}$ restricts to a biholomorphism from $U_{v}$ to $D(0,r_{v},K_{v})$ for some $r_{v}>0$.
For each $v\in\Sigma_{K}$, define
$R^{\leq}_{v}:=\min\{R_{v},r_{v}\}\leq r_{v}$
and $\Delta_{v}:=(x^{v\textrm{-}\mathrm{an}}|_{U_{v}})^{-1}(D(0,R^{\leq}_{v},K_{v}))$. By shrinking the $r_{v}$ (if necessary), we can and do assume that $\Delta_{v}$ is
  • (a)
    simply connected for $v\in\Sigma_{v,\infty}$;
  • (b)
    contained in the open subspace $\mathcal{C}_{v}$ of Section 3.2.2 for the (finitely many) $v\in\Sigma_{v,f}$ of bad reduction.
By Lemma 9, we can and do also suppose that $\Delta_{v}$ is
  • (c)
    simple for all $v\in\Sigma_{K,f}$ of good reduction.
Lemma 10.
[…]
$\mathfrak{G}^{\prime}_{1,\mathfrak{s},k_{\hat{v}}}\cong\mathfrak{G}^{\prime}_{1,\mathfrak{s}_{0},k_{\hat{v}}}\text{ and }\mathfrak{G}^{\prime}_{2,\mathfrak{s},k_{\hat{v}}}\cong\mathfrak{G}^{\prime}_{2,\mathfrak{s}_{0},k_{\hat{v}}}$.
@@
[…]
Now we state the main result of this section.
Proposition 11.
[…]
Let $v\in\Sigma_{K}$ and choose $\hat{v}\in\Sigma_{\hat{K}}$ lying over $v$.
If $s^{\hat{v}\textrm{-}\mathrm{an}}\in\Delta_{v}$, there exists a $\hat{v}-$adic relation $Q_{\hat{v}}$ of degree at most $2$ between the evaluations of the elements of $\mathcal{G}$ at $x(s)$ which is not contained in the ideal $I$ of $\overline{\mathbb{Q}}[X_{imn}:1\leq i,m,n\leq 2]$ generated by the elements
@@
3.2. Proof of Proposition 11
If necessary, replace $K$ with $\hat{K}$ and $v$ with $\hat{v}$, and, for uniformity of notation, define the field
@@ 3.2. Proof of Proposition 11
(This is the field of coefficients for the “$v$-adic cohomology” theory i.e. Betti cohomology when $v\in\Sigma_{K,\infty}$ or crystalline cohomology when $v\in\Sigma_{K,f}$.) We will split the proof of Proposition 3.2 into two cases.

### v2

@@ 2.6. G-functions associated with an abelian scheme
Let $K$, $C$, $s_{0}$, $x$ be as in Section 2.4 and let $\pi\colon\mathcal{A}\to C$ be an abelian scheme of relative dimension $g$, also defined over $K$. We write $H^{1}_{DR}(\mathcal{A}/C)$ for the relative algebraic de Rham cohomology, equipped with its Gauss–Manin connection
@@
By [Del70, Prop. 6.14], the $\mathcal{O}_{C}$-sheaf $H^{1}_{DR}(\mathcal{A}/C)$ is locally-free. Suppose that it is free (which we will later achieve by replacing $C$ with a Zariski open subset containing $s_{0}$). Equivalently, we can choose a basis $\{\omega_{1},\ldots,\omega_{2g}\}$ of global sections of $H^{1}_{DR}(\mathcal{A}/C)$.
Lemma 2.1.
Let $v\in\Sigma_{K}$. There exists an open neighbourhood $\Delta_{v}$ of $s_{0}^{v\textrm{-}\mathrm{an}}$ such that $H^{1}_{DR}(\mathcal{A}/C)^{v\textrm{-}\mathrm{an}}(\Delta_{v})$ possesses a basis of horizontal sections $\{\gamma_{v,1},\ldots,\gamma_{v,2g}\}$.
Proof.
For $v\in\Sigma_{K,\infty}$, the existence of such a $\Delta_{v}$ is classical and follows from Cauchy’s theorem—any simply connected open subset would suffice. For $v\in\Sigma_{K,f}$, the existence of such a $\Delta_{v}$ is guaranteed by [Lut37, Thm. IV]. ∎
Let $v\in\Sigma_{K}$. Let $\Delta_{v}$ and $\{\gamma_{v,1},\ldots,\gamma_{v,2g}\}$ be as in Lemma 2.1. Let
@@
Lemma 2.3.
Define
$Y_{v}:=\Omega_{v}\cdot\Omega_{v}(s^{v\textrm{-}\mathrm{an}}_{0})^{-1}\in\mathrm{M}_{2g}(\mathcal{O}_{C^{v\textrm{-}\mathrm{an}}}(\Delta_{v}))$
and denote its entries by $Y_{v,mn}$ (where $1\leq m,n\leq 2g$). Let $F_{v,mn}$ denote the Taylor series of $Y_{v,mn}$ with respect to $x$. Then the $F_{v,mn}$ are G-functions, belonging to K[[X]]K[\![X]\!], and are independent of $v\in\Sigma_{K}$.
Note that $Y_{v}$ and the $F_{v,mn}$ are independent of the choice of $\gamma_{v,1},\dotsc,\gamma_{v,2g}$. We make use of this in Section 3.2, where we impose additional conditions on this basis and rely on the fact that this leaves the $F_{v,mn}$ unaffected.
@@
It is straightforward to verify that the matrix $F_{v}:=(F_{v,mn})_{mn}$ satisfies a linear differential equation $\frac{d}{dx}(-)=A(-)$ induced from the action of $\nabla(\frac{d}{dx})$ on $H^{1}_{DR}(\mathcal{A}/C)$, where $A\in\mathrm{M}_{2g}(K((X)))$. Starting from the fact that $F_{v}(0)=I_{2g}$ has entries in $K$, it follows that $F_{v,nm}\in K[\![X]\!]$ for all $m,n\in\{1,\ldots,2g\}$. By [And89, p. 3, Thm. B], therefore, $F_{v,nm}$ is a G-function for all $m,n\in\{1,\ldots,2g\}$. Furthermore, $F_{v}$ is the unique solution of $\frac{d}{dx}(-)=A(-)$ in $\mathrm{M}_{2g}(K((X)))$ satisfying $F_{v}(0)=I_{2g}$, so it is independent of $v$. ∎
Because the $F_{v,mn}$ in Lemma 2.3 are independent of $v$, we denote them simply by $F_{mn}$. We refer to the $F_{mn}$ for $1\leq m,n\leq 2g$ as the G-functions associated with $\mathcal{A}\to C$, $x$, $s_{0}$ and $\{\omega_{1},\ldots,\omega_{2g}\}$.
@@ 2.7. Relations
Let $F_{1},\dotsc,F_{n}\in\overline{\mathbb{Q}}[\![X]\!]$. We say that a polynomial $\tilde{Q}\in\overline{\mathbb{Q}}[X][X_{1},\dotsc,X_{n}]$ is a functional relation between $F_{1},\dotsc,F_{n}$ if it is homogeneous in $X_{1},\dotsc,X_{n}$ and
@@
Let $\xi\in\overline{\mathbb{Q}}$ and let $Q\in\overline{\mathbb{Q}}[X_{1},\dotsc,X_{n}]$ be a homogeneous polynomial. Let $K$ be a number field which contains $\xi$ and all of the coefficients of $F_{1},\dotsc,F_{n}$ and $Q$. Following [And89, Ch. VII, 5.1], we make the following definitions:
@@
    $Q$ is a trivial relation between $F_{1},\dotsc,F_{n}$ at $\xi$ if it is the specialisation at $X=\xi$ of a functional relation $\tilde{Q}\in\overline{\mathbb{Q}}[X][X_{1},\dotsc,X_{n}]$ between $F_{1},\dotsc,F_{n}$.
@@ 2.8. Simple neighbourhoods
Return to the situation of Section 2.4 and let $\mathfrak{C}$ be a regular integral flat $\mathcal{O}_{K}$-scheme with an isomorphism $\mathfrak{C}_{K}\to C$. Let $v\in\Sigma_{K,f}$ and let $\hat{\mathfrak{C}}_{\mathfrak{C}_{k_{v}}}$ denote the formal completion of $\mathfrak{C}$ along the fibre $\mathfrak{C}_{k_{v}}$. Let $\hat{\mathfrak{C}}_{\mathfrak{C}_{k_{v}}}^{\mathrm{rig}}$ denote the associated rigid analytic space (see [Ber96, 0.2.6]). Then there is a canonical open immersion $\hat{\mathfrak{C}}_{\mathfrak{C}_{k_{v}}}^{\mathrm{rig}}\to C^{v\textrm{-}\mathrm{an}}$ [Ber96, Prop. 0.3.5] and reduction map $\operatorname{red}_{v}:\hat{\mathfrak{C}}_{\mathfrak{C}_{k_{v}}}^{\mathrm{rig}}\to\mathfrak{C}_{k_{v}}$. We say that an open subspace $U$ of $C^{v\textrm{-}\mathrm{an}}$ is simple (with respect to the integral model $\mathfrak{C}$) if all of the following hold:
  1. (1)
    $U\subset\hat{\mathfrak{C}}_{\mathfrak{C}_{k_{v}}}^{\mathrm{rig}}$;
  2. (2)
    ${\operatorname{red}_{v}}|_{U}$ is constant;
  3. (3)
    the value of $\operatorname{red}_{v}$ on $U$ is a smooth closed point $t\in\mathfrak{C}_{k_{v}}$.
(In other words, $U$ is contained in the residue disc of a smooth closed point of $\mathfrak{C}_{k_{v}}$.)
@@ 3. Relations for elliptic schemes
Let $\mathcal{E}_{1},\mathcal{E}_{2}$ be elliptic schemes over the same curve $C$ having isogenous fibres at some base point $s_{0}$. In this section, we construct $v$-adic relations between the G-functions associated with $\mathcal{E}_{1},\mathcal{E}_{2}$ (over the same curve $C$) at points $s\in C(\overline{\mathbb{Q}})$ for which the fibres are again isogenous. In Sections 3.1 and 3.2, we consider a pair of elliptic schemes, as just described, and construct relations which are potentially different at each place $v$. In Section 3.3, we introduce a third elliptic scheme, in order to obtain a relation independent of $v$, valid at all places $v$ of good ordinary reduction of the fibres over the base point.
@@ 3.1. Isogenous pairs
Let $C$ be a smooth projective geometrically irreducible algebraic curve, defined over a number field $K$, and let $\mathfrak{C}$ be a regular $\mathcal{O}_{K}$-model of $C$ (see [Liu06, Definition 10.1.1] for the definition; such a model exists by [Liu06, Cor. 8.3.45]).
@@
For any finite extension $L$ of $K$ and $s\in C(L)$, we refer to the unique section $\mathfrak{s}\colon\operatorname{Spec}(\mathcal{O}_{L})\to\mathfrak{C}_{\mathcal{O}_{L}}$ whose image is the Zariski closure of $s$ as the $\mathcal{O}_{L}$-extension of $s$ (see [DO23, Lemma 6.3] for the existence of such a section). Note that, for any $v\in\Sigma_{K,f}$, the point $\mathfrak{s}(\mathfrak{p}_{v})\in\mathfrak{C}_{k_{v}}$ is smooth (see [Liu06, Exercise 4.3.25(c)]).
For $i\in\{1,2\}$, let $\mathfrak{G}_{i}\to\mathfrak{C}$ be a semiabelian scheme such that, for some non-empty Zariski open subset $C_{0}\subset C$, the base-change $\mathcal{E}_{i}:=\mathfrak{G}_{i}\times_{\mathfrak{C}}C_{0}$ is an elliptic scheme over $C_{0}$. Let $\mathfrak{C}_{0}\subset\mathfrak{C}$ denote a non-empty Zariski open subscheme such that $\mathfrak{G}_{i,0}:=\mathfrak{G}_{i}\times_{\mathfrak{C}}\mathfrak{C}_{0}$ is an elliptic scheme over $\mathfrak{C}$ for $i\in\{1,2\}$ (such a subscheme exists because the toric rank is upper semi-continuous—see [Lan13, Lemma 3.3.1.4]). Suppose that $\mathfrak{C}_{0,K}=C_{0}$.
Let $s_{0}\in C_{0}(K)$ and suppose that there exists a $K$-isogeny $f_{0}:\mathcal{E}_{1,s_{0}}\to\mathcal{E}_{2,s_{0}}$ of degree $N_{0}:=\deg(f_{0})$. We will say that $v\in\Sigma_{K,f}$ is a place of good (resp. bad) reduction if one (equivalently, both) of the $\mathcal{E}_{i,s_{0}}$ have good (resp. bad) reduction at $v$.
Choose bases $\{\omega_{i},\eta_{i}\}$ for $H^{1}_{DR}(\mathcal{E}_{i}/C_{0})$ such that
(3) $\omega_{1,s_{0}}=\mu f^{*}_{0}\omega_{2,s_{0}}\text{ and }\eta_{1,s_{0}}=\mu f^{*}_{0}\eta_{2,s_{0}},$
for some $\mu\in\mathbb{Q}$, where
@@
denotes the pullback induced by $f_{0}$ ($\mu$ will afford us a degree of freedom which we will use in Section 4). Let $x\in K(C_{0})$ be a local parameter at $s_{0}$ and let $F_{imn}\in K[\![X]\!]$ denote the G-functions associated with $\mathcal{E}_{i}\to C_{0}$, $x$, $s_{0}$ and $\{\omega_{i},\eta_{i}\}$. Let
$\mathcal{G}:=\{F_{imn}:1\leq i,m,n\leq 2\}.$
For each $v\in\Sigma_{K,\infty}$, let $r_{v}$ and $U_{v}$ denote, respectively, a positive real number and an open neighbourhood of $s_{0}^{v\textrm{-}\mathrm{an}}$ in $C^{v\textrm{-}\mathrm{an}}_{0}$ such that $x^{v\textrm{-}\mathrm{an}}$ restricts to an isomorphism of $K_{v}$-analytic spaces from $U_{v}$ to $D(0,r_{v},K_{v})$. For the $v\in\Sigma_{K,f}$, let $r_{v}$ and $U_{v}$ be the positive real numbers and simple open subspaces of $C^{v\textrm{-}\mathrm{an}}_{0}$ afforded to us by Proposition A.2. Define
$R_{v}:=\min\{1,r_{v},R(F^{v\textrm{-}\mathrm{an}}):F\in\mathcal{G}\}\leq r_{v}$
and $\Delta_{v}:=(x^{v\textrm{-}\mathrm{an}}|_{U_{v}})^{-1}(D(0,R_{v},K_{v}))$.
Note that we have $R_{v}=1$ for almost all $v\in\Sigma_{K}$. Moreover, $\Delta_{v}$ is simply connected for $v\in\Sigma_{K,\infty}$ and simple for $v\in\Sigma_{v,f}$.
Lemma 3.1.
Let $s\in C(L)$ for some finite extension $L$ of $K$. Let $v\in\Sigma_{K,f}$ be a place of good reduction and let $w$ be a finite place of $L$ lying above $v$ such that $s^{{w\textrm{-}\mathrm{an}}}\in\Delta_{v}$. Then, if $\mathfrak{s},\mathfrak{s}_{0}:\operatorname{Spec}\mathcal{O}_{L}\to\mathfrak{C}_{\mathcal{O}_{L}}$ denote the $\mathcal{O}_{L}$-extensions of $s$ and $s_{0}$, respectively, there are canonical isomoprhisms
$\mathfrak{G}_{1,\mathfrak{s},k_{w}}\cong\mathfrak{G}_{1,\mathfrak{s}_{0},k_{w}}\text{ and }\mathfrak{G}_{2,\mathfrak{s},k_{w}}\cong\mathfrak{G}_{2,\mathfrak{s}_{0},k_{w}}.$
@@
Since $\Delta_{v}$ is simple, we have $\operatorname{red}_{w}(s^{w\textrm{-}\mathrm{an}})=\operatorname{red}_{w}(s^{w\textrm{-}\mathrm{an}}_{0})$. Equivalently, $\mathfrak{s}(\mathfrak{p}_{w})=\mathfrak{s}_{0}(\mathfrak{p}_{w})$. The isomorphisms now follow by definition. ∎
We refer to the above data, namely,
$\{C,\mathfrak{C},\mathfrak{G}_{i},C_{0},\mathcal{E}_{i},\mathfrak{C}_{0},s_{0},f_{0},\{\omega_{i},\eta_{i}\},x,F_{imn},R_{v},\Delta_{v}:1\leq i,m,n\leq 2,v\in\Sigma_{K}\}$
as an isogeny datum. The main result of this section is as follows.
Proposition 3.2.
Consider an isogeny datum as above.
Let $L$ be a finite extension of $K$ and suppose $s\in C_{0}(L)$ is such that there exists an $L$-isogeny $f:\mathcal{E}_{1,s}\to\mathcal{E}_{2,s}$.
Let $v\in\Sigma_{K}$ and choose $w\in\Sigma_{L}$ lying over $v$. Suppose that $s^{w\textrm{-}\mathrm{an}}\in\Delta_{v}$.
There exists a $w$-adic relation $Q_{w}$ of degree at most $2$ between the evaluations of the elements of $\mathcal{G}$ at $x(s)$ which is not contained in the ideal $I$ of $\overline{\mathbb{Q}}[X_{imn}:1\leq i,m,n\leq 2]$ generated by the elements
@@
3.2. Proof of Proposition 3.2
If necessary, replace $K$ with $L$ and $v$ with $w$ and, for uniformity of notation, define the field
@@
(This is the field of coefficients for the “$v$-adic cohomology” theory, namely, Betti cohomology when $v\in\Sigma_{K,\infty}$ or crystalline cohomology when $v\in\Sigma_{K,f}$.) We will split the proof of Proposition 3.2 into two cases.